# Prethermalization in Fermi-Pasta-Ulam-Tsingou chains

### v1

@@ Abstract
The observation of the Fermi-Pasta-Ulam-Tsingou (FPUT) paradox, namely the lack of equipartition in the evolution of a normal mode in a nonlinear chain on unexpectedly long times, is arguably the most famous numerical experiment in the history of physics. Since seventy years after its publication, most studies in FPUT chains still focus on long wavelength initial states similar to the original paper. It is shown here that all characteristic features of the FPUT paradox are rendered even more striking if short(er) wavelength modes are evolved instead. Since not every normal mode leads to equipartition, we also provide a simple technique to predict which modes, and in what order, are excited starting from an initial mode (root) in $\alpha$-FPUT chains. The excitation sequences associated with a root are then shown to spread energy at different speeds, leading to prethermalization regimes that become longer as a function of mode excitation number. This effect is visible in observables such as mode energies and spectral entropies and, surprisingly, also in the time evolution of invariant quantities such as Lyapunov times and Kolmogorov-Sinai entropies. Our findings generalize the original FPUT experiment, provide an original look at the paradox’s source, and enrich the vast literature dedicated to studying equipartition in classical many-body systems.
@@ I Introduction
The numerical experiment of Fermi, Pasta, Ulam and Tsingou (FPUT) [1] was simultaneously the first computational approach in the study of out-of-equilibrium many-body systems and a source of considerable headache shared by generations of physicists in the past 70 years [1, 2, 3, 4]. A proper understanding of FPUT’s findings, which contradicted standard knowledge at the time and were deemed a paradox, relies heavily on modern mathematical and physical tools unavailable to the authors at that time, e.g. a mature formulation of Kolmogorov-Arnol’d-Moser (KAM) theory [5, 6, 7], resonance overlap criteria [8] and Birkhoff-Gustavson normalization [9, 10, 11, 12, 13], and also on the power of modern computers. Being at the intersection of several fields, it is not surprising that this topic has attracted the attention of so many researchers working on nonlinear dynamical systems and statistical mechanics.
The original FPUT experiment consists in evolving the fundamental mode of a harmonic chain in a weakly perturbed, nonintegrable system of which this mode is no longer a stationary state. More specifically, the authors studied the evolution of the first normal mode of a discretized string under the dynamics generated by the Hamiltonian
@@
where $p_{j}$ and $x_{j}$ are the momentum and position of each discretized mass in the chain, $\alpha{=}1/4$ and fixed boundary conditions were used, namely $x_{0}{=}x_{N+1}=0$. A very small energy density was chosen for the initial state, i.e. $h{=}H/N\ll 1$, limiting the size of the cubic term and rendering the system a very weak perturbation of the integrable harmonic chain corresponding to setting $\alpha{=}0$. Even for such a weak perturbation the dynamics under (1) ends up being chaotic, justifying the original FPUT expectation that the energy of the initially excited mode would be shared among all other normal modes in a finite (possibly short) time. However, what was verified is that not only energy equipartition was not reached for very long times, but that the energy periodically returned to the initial state and formed recurrences, with a large fraction of modes being barely excited during evolution.
@@
Nowadays, the intersection of a plethora of studies provides a rather converged explanation for the originally observed apparent lack of energy equipartition in finite, weakly anharmonic systems (for reviews, see [2, 14, 15, 16, 4, 17]), although speculations on whether or not the original FPUT setup would eventually reach equilibrium have become a prominent topic in metaphysics. First of all, it is important to note that the energy density chosen by FPUT was extremely small; indeed, a subsequent investigation by Chirikov and Izrailev showed that larger energy densities did result in energy equipartition [18]. Secondly, KAM theory was later proved to be applicable to many types of FPUT-like systems, showing that there is indeed a bound in energy density below which equipartition will never take place [19, 20, 21, 22]. Thirdly, and most relevant to this manuscript, is the fact that the initial state chosen by FPUT is a highly atypical initial state, being a stationary solution of a nearby harmonic chain [23]. Even more, such an atypical initial state is also close to exact stationary time-periodic states coined q-breathers of the anharmonic system under consideration [24, 25]. This strongly contrasts with typical initial conditions, which when expressed in the normal mode basis are combinations of all of its elements. Since such modes form the canonical action variables of the harmonic chain, perturbing a typical initial condition involves an instantaneous coupling of all actions, while for the atypical FPUT initial state only a single action is perturbed. It is therefore only logical that time evolution would be atypical in the original FPUT study, given the atypicality already present in its chosen initial state.
@@
Atypical initial states are, moreover, not all built the same. Their time evolution behaves very differently depending on which initial mode, or root, is excited. The dynamics of roots is characterized by excitation sequences that spread in mode space and, depending on relatively simple algebraic rules, completely or only partially fill it. The case of complete filling is associated to thermal roots, namely states for which time-dependent observables will generally reach equipartition; If the root only excites a subset of modes, known as its bush [26], then equipartition does not take place and the corresponding root is deemed nonthermal. A particularly striking example of nonthermal roots is given by q-breathers, namely families of exact time-periodic orbits which localize in mode space and reduce to the root mode upon tuning the anharmonicity down to zero [24, 25]. Such states show that specific normal modes and their perturbations can also be stationary states of the perturbed chain and provide remarkable counterexamples to the expectation that observables in nonlinear, classical many-body systems necessarily undergo some form of equilibration. They also show that ergodicity-breaking behavior can be observed not only as a function of decreasing energy density, as in the original FPUT experiment, but also as a function of the initial state, as we shall explore here.
@@
This manuscript is devoted to investigating the properties of atypical initial conditions (normal modes) in FPUT chains, and also comparing them to the ones of typical (random) ones. We provide explicit algebraic rules to write mode excitation sequences as a function of the initial root, together with simple criteria to know if such root is thermal or not. We then numerically investigate the time evolution of thermal roots of several different excitation numbers, showing that the paradox associated to the fundamental mode is substantially maximized for higher excited roots, e.g. the prethermal regime and its metastable plateaus become increasingly longer as a function of root excitation number for a fixed energy density. This is shown both by a direct tracking of mode energies as a function of time and by computing equipartition times from spectral entropies.
@@
The impact of root excitation number is not restricted to quantities that depend on the initial state. Indeed, albeit Lyapunov times and Kolmogorov-Sinai entropies being invariant with respect to a large number of transformations and being constant in an ergodic system regardless of the initial state used for computing them [27], we show here that their transient properties are strikingly different from those of typical initial states and clearly reflect the fact that the system is trapped in phase space. More specifically, both the maximal Lyapunov exponent (the inverse of the Lyapunov time) and the Kolmogorov-Sinai entropy show a dip in their time evolution when computed from roots, which is absent if the initial state chosen is typical. Moreover, the duration and depth of this dip increases with excitation number, showing that excited modes remain trapped for even longer in near-integrable portions of phase space. An explanation to this phenomenon is provided by the previously computed excitation sequences of roots, which show that equipartition is fastest for states which spread energy efficiently among modes. The fastest spreader in this sense is always the fundamental root, with the equipartition time slowly increasing with excitation number. Evidently, nonthermal roots are extremely bad spreaders since a portion of mode space is never filled.
@@ II Model and observables
We restrict our analysis to cubic FPUT chains, also known as the $\alpha$-FPUT model, with Hamiltonian (1). In the following we recall how to rewrite the FPUT Hamiltonian in normal mode (phonon) basis, and introduce quantities such as mode energies, spectral entropies and equipartition times. We then describe how to obtain important invariant quantities that will be used in the following sections, such as Lyapunov times and Kolmogorov-Sinai entropies.
II.1 Eigenmode basis
Since (1) has $N$ degrees of freedom its harmonic limit has a set of $N$ eigenmodes, denoted $\{P_{j},Q_{j}\}_{j}$, with which we can rewrite the position and momentum in (1) at any time $t$ as
@@ II.1 Eigenmode basis
each mode having eigenfrequency and eigenenergy
@@ II.2 Spectral entropy
Eq. (1) is much easier to deal with, numerically speaking, than (II.1). However, we are interested in several properties that need to be computed in the eigenmode basis, such as the eigenenergies themselves and the [normalized] spectral entropy
@@
where the Shannon entropy and normalized eigenenergy are given, respectively, by
@@
Spectral entropy has been used several times in order to quantify equipartition in nonlinear systems [23, 4, 29, 30]. The logic for employing it is simple: If ideal equipartition is reached, $\rho_{j}(t)$ will be stationary and the same for every mode $j$, freezing $\eta(t)$ at some value $\langle\eta\rangle$. If one assumes a Gibbs distribution at equilibrium, $\langle\eta\rangle$ can be calculated analytically as a function of the Shannon entropy of the initial state, being given by
@@ II.3 Lyapunov invariants and Kolmogorov-Sinai entropy
Equipartition time as defined in Sec. II.2 is, evidently, attached to the observables being monitored, namely the mode energies. Different observables will generally reach equipartition at different times (if at all), such that observable equipartition does not actually measure properties intrinsic to the system at hand. Nevertheless, since mode energies are conserved in the unperturbed harmonic chain, one knows well how these observables should behave upon approaching an integrable limit, i.e., equipartition should slow down and eventually stop for sufficiently small energy densities.
There is, however, an observable-independent time scale that is completely intrinsic to the system and should not depend on the initial state, given by the Lyapunov time (the inverse of the maximal Lyapunov exponent). Since, like the maximal Lyapunov exponent (MLE), all exponents in the Lyapunov spectrum (LS) are invariant with respect to homeomorphisms [27], the LS provides a valuable set of time scales related to the average “pull” along each 1-dimensional submanifold that forms the hyperbolic skeleton of the system’s dynamics in phase space. The information in the LS can also be condensed in the form of a scalar, the Kolmogorov-Sinai entropy, which by Pesin’s theorem can be obtained as the sum of all positive exponents in the LS [32].
In order to compute the LS we employ the well-known prescription of [33, 34], which amounts to computing the spectrum of the symmetric matrix
@@
In the above $\mathcal{M}$ represents the system’s stability matrix, which is obtained by solving Hamilton’s equations in tangent space:
@@
where $\mathrm{Hess}$ is the Hessian with respect to $(\boldsymbol{p},\boldsymbol{x})$. Since the stability matrix has to be computed along a trajectory, the above equation is solved in parallel with Hamilton’s equations for an initial phase-space point $(\boldsymbol{p}_{0},\boldsymbol{x}_{0})$ until a final time that is long enough to result in a converged LS. During this evolution, one must also perform QR-diagonalizations of $\mathcal{M}$ in order to avoid numerical denegenacies, as is well-known and described in, e.g., [34].
As can be seen in (9), the stability matrix will generally depend on the trajectory along which it is calculated, i.e., on the initial phase-space point $(\boldsymbol{p}_{0},\boldsymbol{x}_{0})$. However, for a sufficiently large number of degrees of freedom and energy density, phase space will consist of a single metrically intransitive set that is equally accessible to all trajectories in the system [35, 36]. Equivalently, one can formulate this previous property as stating that, for sufficiently large $N$ and $h{=}E/N$, the system lies outside the KAM regime in which phase space can be split into a chaotic web and pockets of near-integrability. Thus, since all trajectories are “free to roam” around the entire accessible phase space, the LS is expected to be independent of the trajectory along which it is computed. This independence is a stronger evidence that the system at hand is ergodic than computing observables, and can be ascertained directly from the verification that the KS entropy is independent of the trajectory along which it is computed.
@@ III Excitation sequences
We start this section by writing a very abridged version of Newton’s second law for a given mode $i$, obtained from (II.1):
$\ddot{Q}_{i}+\omega_{i}^{2}Q_{i}\sim\sum_{j,k=1}^{N}B_{i,j,k}\,x_{j}Q_{k}\,.$ (11)
Thus, mode $Q_{i}$ is excited either by a lone mode $x_{j}$ or a mixture of modes $x_{j}$ and $Q_{k}$. Excitations from single modes have $j=k$ and the coupling tensor assumes the form
@@
The above expression is composed of: An ascending term, namely $\delta_{i,k+j}$, which dictates which higher-order mode to excite starting from $j$ and $k$; A right reflection term, $\delta_{i,2(N+1)-(j+k)}$, which dictates which term to excite if $j+k>(N+1)$, i.e. if the next mode in the ascending sequence is too large and leaves mode space; And a descending term, $\delta_{i,|k-j|}$, which propagates the excitation backward in mode space and is also responsible for controlling left reflections, which take place when the next excited mode in a descending sequence is smaller than zero.
III.1 First excitation cycle and q-breathers
Initializing the system in a root $Q_{r}$ means setting $x_{j}{=}0$ unless $j{=}r$. The first excitation cycle, therefore, starts with a lone mode $Q_{r}$, which by (III) excites modes $i=2r$ and $i=2(N-r+1)$. If the first condition is met, then the second delta will be one iff $r=(N+1)/2$, with $N+1$ even. Thus, if the chain has an odd number of sites and the root chosen is $Q_{(N+1)/2}$, then all coefficients in the coupling tensor vanish and the energy remains forever localized in the root, i.e. this mode is a q-breather.
Now, assuming that the second Kronecker delta in (III) vanishes, i.e. that $i\neq 2(N-r+1)$, then the only mode that is excited starting from $Q_{r}$ is $Q_{2r}$. This constitutes the first excitation cycle, namely the first run over $j$ and $k$ indices in (11). Note that $Q_{r}$ excites $Q_{2r}$ only if $2(r-1)<N$, since $i$ cannot be negative or zero. This justifies splitting excitation sequences into two types, namely low- and high-mode excitations, depending on whether the root $r$ fulfills $2(r-1)<N$ or not. The only difference is that for low-mode excitations the initial sequence will be ascending, while it is descending in the high-mode case. Although these scenarios are essentially identical, we will assume low-mode excitations here in order for sequences to start in the ascending phase.

### v2

@@ Abstract
The observation of the Fermi-Pasta-Ulam-Tsingou (FPUT) paradox, namely the lack of equipartition in the evolution of a normal mode in a nonlinear chain on unexpectedly long times, is arguably the most famous numerical experiment in the history of physics. Seventy years after the original publication, most studies in FPUT chains still focus on long wavelength initial states similar to the original paper. It is shown here that all characteristic features of the FPUT paradox are rendered even more striking if modes with short(er) wavelengths are evolved instead. Since not every normal mode leads to equipartition, we also provide a simple technique to predict which modes, and in what perturbation order, are excited starting from an initial mode (root) in $\alpha$-FPUT chains. The excitation sequences associated with a root are then numerically shown to spread energy at different speeds, leading to prethermalization regimes that become longer as a function of mode excitation number. This effect is visible in observables such as mode energies and spectral entropies and, surprisingly, also in the time evolution of invariant quantities such as Lyapunov times and Kolmogorov-Sinai entropies. Our findings generalize the original FPUT experiment, provide an original look at the paradox’s source, and enrich the vast literature dedicated to studying equipartition in classical many-body systems.
@@ I Introduction
The numerical experiment of Fermi, Pasta, Ulam and Tsingou (FPUT) [1] was simultaneously the first computational approach in the study of out-of-equilibrium many-body systems and a source of considerable headache shared by generations of physicists in the past 70 years [1, 2, 3, 4]. A proper understanding of FPUT’s findings–which contradicted the standard knowledge of the time and were considered paradoxical–relies heavily on modern mathematical and physical tools unavailable to the original authors, e.g., a mature formulation of Kolmogorov-Arnol’d-Moser (KAM) theory [5, 6, 7], resonance overlap criteria [8] and Birkhoff-Gustavson normalization [9, 10, 11, 12, 13], and also on the power of modern computers. Sitting at the intersection of several fields, it is no surprise that this topic has drawn sustained interest from researchers in nonlinear dynamical systems and statistical mechanics.
The original FPUT experiment consists of evolving the fundamental mode of a harmonic chain in a weakly perturbed, nonintegrable system in which this mode is no longer a stationary state. More specifically, the authors studied the evolution of the first normal mode of a discretized string governed by the Hamiltonian
@@
where $p_{j}$ and $x_{j}$ are the momentum and position of each discretized mass in the chain, $\alpha{=}1/4$ and fixed boundary conditions were used, namely $x_{0}{=}x_{N+1}=0$. A very small energy density was chosen for the initial state, i.e., $h{=}H(\boldsymbol{p},\boldsymbol{x})/N\ll 1$, which limited the contribution of the cubic term and rendered the system a weak perturbation of the integrable harmonic chain obtained by setting $\alpha{=}0$. Even with such a weak perturbation, the dynamics under (1) is chaotic, which motivated the original expectation that the energy of the initially excited mode would rapidly spread among all other normal modes, leading to equipartition. However, what was observed was quite different: Not only did energy equipartition fail to occur over long timescales, but the system exhibited near recurrences, with energy periodically returning to the initial mode and many other modes remaining only weakly excited throughout the evolution [14].
@@
Today, the convergence of numerous studies provides a coherent explanation for the originally observed lack of energy equipartition in finite, weakly anharmonic systems (for reviews, see [2, 15, 16, 17, 4, 18]) – although speculations on whether or not the original FPUT setup would eventually reach equilibrium have become a prominent topic in metaphysics. First, it is important to note that the energy density chosen by FPUT was extremely small; indeed, a subsequent investigation by Chirikov and Izrailev showed that larger energy densities do lead to energy equipartition [19]. Second, KAM theory was later shown to apply to many FPUT-like systems, rigorously proving the existence of an energy threshold below which equipartition does not take place [20, 21, 22, 23]. Third, and most relevant to this manuscript, is the fact that the initial condition used by FPUT is highly atypical, corresponding to a stationary solution of the nearby harmonic chain [24]. Moreover, this atypical initial state lies close to exact, time-periodic stationary solutions of the anharmonic system, known as q-breathers [25, 26]. This stands in sharp contrast to typical initial conditions which, when expressed in the normal mode basis, involve combinations of all of its elements. Since these modes form the canonical action variables of the harmonic chain, perturbing a typical initial condition leads to an immediate coupling of all actions, whereas in the FPUT case, only a single action is perturbed. It is therefore natural that the resulting time evolution in the original FPUT study was highly atypical, reflecting the atypicality already present in the initial state itself.
A further important point is that atypical initial states are not all alike. Their time evolution depends sensitively on which initial mode, or root, is excited. The dynamics of roots is characterized by excitation sequences that spread in mode space and, depending on relatively simple algebraic rules, completely or only partially fill it. The case of complete filling corresponds to thermal roots, for which time-dependent observables generally reach equipartition. In contrast, if the root excites only a subset of modes, known as its bush [27], then equipartition fails to occur, and the corresponding root is termed nonthermal.
@@
A particularly striking example of nonthermal roots is provided by q-breathers: families of exact, time-periodic solutions that remain localized in mode space and reduce to a single-mode excitation as the anharmonicity tends to zero [25, 26]. These states show that certain normal modes and their perturbations can remain stationary even in the presence of nonlinearity, offering remarkable counterexamples to the expectation that observables in nonlinear classical many-body systems necessarily undergo some form of equilibration. They also reveal that ergodicity-breaking behavior can arise not only as a function of decreasing energy density, as in the original FPUT experiment, but also as a consequence of the choice of initial state, as we shall explore in detail here.
@@
This manuscript is devoted to investigating the properties of atypical initial conditions (normal modes) in FPUT chains, and also comparing them to the ones of typical (random) ones. We provide explicit algebraic rules to determine which modes are excited given an initial root, together with their short-time perturbation order in $\alpha$. We then numerically investigate the time evolution of thermal roots of several different excitation numbers, showing that the paradox associated to the fundamental mode is substantially maximized for higher excited roots, e.g., the prethermal regime and its metastable plateaus become increasingly longer as a function of root excitation number for a fixed energy density. This is shown both by a direct tracking of mode energies as a function of time and by computing equipartition times from spectral entropies.
@@
We also show that the root excitation number impacts quantities that do not depend on the initial state. Indeed, albeit Lyapunov times and Kolmogorov-Sinai (KS) entropies being invariant with respect to a large number of transformations and being constant in an ergodic system regardless of the initial state used for computing them [28], we show here that their transient properties are strikingly different from those of typical initial states and clearly reflect the fact that the system is trapped in phase space [24]. More specifically, both the maximal Lyapunov exponent (the inverse of the Lyapunov time) and the Kolmogorov-Sinai entropy show a dip in their time evolution when computed from roots, which is absent if the initial state chosen is typical. Moreover, the duration and depth of this dip increases with excitation number, showing that excited modes remain trapped for even longer in near-integrable portions of phase space, likely due to their proximity to q-breathers.
@@ II Model and observables
We restrict our analysis to cubic FPUT chains, also known as the $\alpha$-FPUT model, with Hamiltonian (1). In the following we recall how to rewrite the FPUT Hamiltonian in normal mode (or phonon, or eigenmode) basis, and introduce quantities such as mode energies, spectral entropies and equipartition times. We then describe how to obtain important invariant quantities that will be used in the following sections, such as Lyapunov times and Kolmogorov-Sinai entropies.
II.1 Normal mode basis
Since Eq. (1) has $N$ degrees of freedom its harmonic limit has a set of $N$ normal modes, denoted $\{P_{j},Q_{j}\}_{j}$, with which we can rewrite the position and momentum in Eq. (1) at any time $t$ as
@@ II.1 Normal mode basis
each mode having frequency and energy
@@ II.2 Spectral entropy
Equation (1) is much easier to numerically deal with than (II.1). However, we are interested in several properties that need to be computed in the normal mode basis, such as the mode energies themselves and the [normalized] spectral entropy
@@
where the Shannon entropy and normalized energies are given, respectively, by
@@
Spectral entropy has been used several times to quantify equipartition in nonlinear systems [24, 4, 30, 31]. The logic for employing it is simple: If equipartition is reached, then $\rho_{j}(t)$ will be stationary and the same for every mode $j$, freezing $\eta(t)$ at some value $\langle\eta\rangle$. If one assumes a Gibbs distribution at equilibrium, then $\langle\eta\rangle$ can be calculated analytically as a function of the Shannon entropy of the initial state and is given by
@@ II.3 Lyapunov invariants and Kolmogorov-Sinai entropy
Equipartition time as defined in Sec. II.2 is, evidently, attached to the observables being monitored, namely the mode energies. Different observables will generally reach equipartition at different times (if at all), such that observable equipartition does not actually measure properties intrinsic to the system. Nevertheless, since mode energies are conserved in the unperturbed harmonic chain, one knows well how these observables should behave upon approaching an integrable limit, i.e., equipartition should slow down and eventually stop for sufficiently small energy densities.
There is, however, an observable-independent timescale that is completely intrinsic to the system and should not depend on the initial state, given by the Lyapunov time (the inverse of the maximal Lyapunov exponent). Since, like the maximal Lyapunov exponent (MLE), all exponents in the Lyapunov spectrum (LS) are invariant with respect to homeomorphisms [28], the LS provides a valuable set of timescales related to the average “pull” along each 1-dimensional submanifold that forms the hyperbolic skeleton of the system’s dynamics in phase space [33]. The information in the LS can also be condensed in the form of a scalar, the Kolmogorov-Sinai entropy, which by Pesin’s theorem can be obtained as the sum of all positive exponents in the LS [34].
To compute the LS we employ the well-known prescription of [35, 36], which amounts to computing the spectrum of the symmetric matrix
@@
In the above, $\mathcal{M}$ represents the system’s stability matrix, obtained by solving Hamilton’s equations in tangent space:
@@
where $\mathrm{Hess}$ is the Hessian with respect to $(\boldsymbol{p},\boldsymbol{x})$. Since the stability matrix has to be computed along a trajectory, the above equation is solved in parallel with Hamilton’s equations for an initial phase-space point $(\boldsymbol{p}_{0},\boldsymbol{x}_{0})$ until a final time long enough to result in a converged LS. During this evolution, one must also perform QR-diagonalizations of $\mathcal{M}$ to avoid numerical denegenacies, as is well-known and described in, e.g., Ref. [36].
As can be seen in (9), the stability matrix will generally depend on the trajectory along which it is calculated, i.e., on the initial phase-space point $(\boldsymbol{p}_{0},\boldsymbol{x}_{0})$. However, for a sufficiently large number of degrees of freedom and energy density, phase space will consist of a single metrically intransitive set that is equally accessible to all trajectories in the system [37, 38]. Equivalently, one can formulate this previous property as stating that, for sufficiently large $N$ and $h{=}H(\boldsymbol{p},\boldsymbol{x})/N$, the system lies outside the KAM regime in which phase space can be split into a chaotic web and pockets of near-integrability. Thus, all trajectories are “free to roam” around the entire accessible phase space, and the LS is expected to be independent of the trajectory along which it is computed. This independence is a stronger evidence of ergodicity than any conclusion drawn from time-dependent observables, and can be verified more succinctly by comparing KS entropies instead of full LS.
@@ III Excitation sequences
We start this section by writing an abridged version of Newton’s second law for a given mode $i$, obtained from (II.1):
$\ddot{Q}_{i}+\omega_{i}^{2}Q_{i}\sim\alpha\sum_{j,k=1}^{N}B_{i,j,k}\,Q_{j}Q_{k}\,.$ (11)
Thus, the mode $Q_{i}$ is excited by a single mode $Q_{j}$ or a mixture of modes $Q_{j}$ and $Q_{k}$. Excitations from single modes have $j{=}k$ and the coupling tensor assumes the form
@@
The above expression is composed of an ascending term, namely $\delta_{i,k+j}$, which dictates which higher-order mode to excite starting from $j$ and $k$; a right reflection term, $\delta_{i,2(N{+}1){-}(j{+}k)}$, which dictates which term to excite if $j{+}k{>}(N{+}1)$, i.e., if the next mode is too large and leaves mode space; and a descending term, $\delta_{i,|k-j|}$, which propagates the excitation backward in mode space and is also responsible for controlling left reflections, which take place when the next excited mode is smaller than zero.
III.1 First excited mode and q-breathers
Initializing the system in a root $Q_{r}$ means setting $Q_{j}{=}0$ unless $j{=}r$. The first excitation cycle, therefore, starts with a lone mode $Q_{r}$, which by (III) excites the modes $i{=}2r$ and $i{=}2(N{-}r{+}1)$. If the first condition is met, then the second delta will be one iff $r{=}(N{+}1)/2$, with $N{+}1$ even. Thus, if the chain has an odd number of sites and the chosen root is $Q_{(N+1)/2}$, then all coefficients of the coupling tensor vanish and the energy remains forever localized in the root: this mode is a q-breather.
Now, assuming that only the second Kronecker delta in (III) vanishes, then the only mode that is excited starting from $Q_{r}$ is $Q_{2r}$ [39, 40, 26]. If, however, only the first delta vanishes, then the first excited mode is obtained after a reflection and is given by $Q_{2(N-r+1)}$. These scenarios are fundamentally identical but, for simplifying writing excitation sequences, we will always assume that the root mode fulfills $2(r{-}1){<}N$, so that the first excited mode will always be $Q_{2r}$.
Moving foward, a naïve perturbative argument allows us to estimate how the amplitude of $Q_{2r}$ scales with the anharmonicity parameter, $\alpha$, for short times. It goes as follows: Since $Q_{r}$ carries all the initial energy in the system and $\alpha$ is small, the solution $Q_{r}(t)$ starts close to a sine or a cosine, as can be seen by setting $\alpha{=}0$ in (11). This approximation scales as $\mathcal{O}(\alpha^{0})$ [26]. Now, $Q_{2r}$ emerges from the product $Q_{r}Q_{r}$ in (11), which is also of 0th order in $\alpha$, but the $\alpha$ multiplying the right-hand side of (11) gives $Q_{2r}(t)$ a $\mathcal{O}(\alpha^{1})$ dependence for short times [41].